# Reputation in the Shadow of Exit

### v5

@@ 2.2 The Repeated Game
The long-run player plays the stage game in each of infinitely many periods ($t=0,1,2,\dots$), weighting each period equally. In each period, they face a population of short-run players who take myopic best replies.
Time-$t$ histories are elements $h^{t}=(y_{s},c_{s})_{s=0}^{t-1}$ of past short-run player signals and long-run player actions. Let $H^{t}$ denote the set of time-$t$ histories, $H=\bigcup_{t}H_{t}$ the set of all histories, and $H^{\infty}=(Y\times C)^{\infty}$ the set of all infinite histories. Short-run player strategies are functions $\sigma_{i}:H\to\Delta(A_{i})$, while the (rational) long-run player’s strategy is a function $\sigma_{L}:H\to\Delta(C)$. [^8] Following Mertens et al. (2015), there is a canonical mapping from strategies $\sigma:H\to\Delta(\mathcal{X})$ and strategies $\sigma:H^{\infty}\to\Delta(\mathcal{X})$ for $\mathcal{X}\in\{C,A\}$. When it is useful (i.e. when topologizing $\sigma$), I will use these two codomains interchangeably.
Given a sequence of realized actions $(a_{t},c_{t})_{t=0}^{\infty}$, the long-run player’s payoff is
@@
Note payoffs here are not discounted. SUPPLEMENTAL APPENDIX studies the discounted repeated game and shows the game here is the standard limit as players become patient, but with exit. However, discounting payoffs does not add anything to the analysis and, given the exit structure, would in fact substantially complicate the analysis. Consequently, I omit the discount rate $\delta$ in the main body of the paper, and refer to this as the case of “extreme patience” for the long-run player.
@@
and let $c^{*}$ be a maximizer in $V_{*}$ (throughout, suppose such a maximizer exists). Let $V^{*}$ be the upper Stackelberg payoff, defined analogously.
@@
The long-run player can either be a rational type, $\omega_{R}\in\Omega$, who seeks to maximize their discounted payoff, or one of potentially many committed types, $\omega_{c}\in\Omega$, who plays $c$ at every possible history. Short-run players hold a full-support prior belief $\mu_{0}\in\Delta(\Omega)$ over the long-run player’s type. Write $\sigma(\cdot,\omega_{c})\equiv c$ for the strategy of a commited type $\omega_{c}$.
For any strategy $\sigma$, define an associated belief system $\mu:H\to\Delta(\Omega)$ which outputs some belief about the long-run players’ type for all histories. Together, $(\sigma,\mu)$ is an assessment. Given an assessment, let $\overline{\sigma}_{L}(h^{t})$ denote the unconditional strategy of the long-run player averaging over $\mu(h^{t})$, and $\sigma_{L}(h^{t};\omega)$ denote the strategy of a type $\omega$ long-run player. Analogously, let $\overline{\mathbb{P}}^{\sigma}$ denote the probability measure over histories generated by the unconditional strategy $\sigma$, and $\mathbb{P}^{\sigma}$ the measure generated when $\omega=\omega_{R}$. The main results of this paper study the set of (rational) long-run player payoffs that can be attained in sequential equilibria $(\sigma,\mu)$, defined as follows:
@@
    Sequential Rationality. At every history $h^{t}$, the continuation strategy $\sigma_{L}(\cdot|h^{t})$ maximizes the long-run player’s continuation payoff given $\{\sigma_{i}(h^{t})\}$ and $\mu(\cdot|h^{t})$.
@@
    Consistency. There exists a sequence $(\sigma^{n},\mu^{n})$, with $(\sigma^{n},\mu^{n})\to(\sigma,\mu)$ such that
@@
Because $\sigma:H\to\Delta(C)$ maps an infinite set into another, the notion of convergence is not immediately clear. Say $(\sigma^{n},\mu^{n})$ converges to $(\sigma,\mu)$ if $\sigma^{n}\to\sigma$ converges in the strong topology (see Appendix B of Luo (2025)) and $\mu^{n}\to\mu$ converges weakly. At any history $h^{t}$ where $\mu(h^{t})$ is degenerate, sequential equilibria are equivalent to subgame perfect equilibria. Consistency of equilibrium profiles will be useful when evaluating the value of deviations the long-run player may consider which are off the path of play (and not supported by some commitment type); absent some consistency-type condition, the value of such deviations would be hard to define in a principled way. Note the “trembles” defined here for consistency do not change the prior belief over $\Omega$ (which would then change $\mu^{n}$). A standard compactness argument shows this is without loss of generality (one could add the requirement $\mu_{0}^{n}\to\mu_{0}$ as well), but for simplicity I omit this additional notation.
Markov perfection requires the stronger assumption that at any two histories where short-run player posterior beliefs about long-run play are the same, players take the same actions. [^9] It may seem unnatural to require the long-run player take the same action when short-run player beliefs are the same. The definition of Markov perfection can be relaxed to require only that short-run players strategies are measurable in $\mu$; however, since there is a single long-run player, their strategy can then be taken to be (without loss of generality) also measurable in $\mu$, less a public randomization device. While such a refinement may seem a-priori restrictive, I show it arises naturally when studying the effect exit has on reputational incentives (see Theorem 1).
Let $\underline{U}$ and $\overline{U}$ be the long-run player’s lowest and highest payoff from any sequential equilibrium in the game. [^10] There is some potential ambiguity about whether sequential equilibrium need to exist when $C$ is infinite (see Myerson and Reny (2020) for some discussion and general fixes). However, given $\Omega$ is finite and $C$ is countable, the problem is substantially simpler—I show in Proposition 1 and Theorem 1 that sequential equilibria as defined above exist, so in particular $\underline{U},\overline{U}$ are finite. Similarly, let $\underline{U}^{M}$ and $\overline{U}^{M}$ be the long-run player’s lowest and highest payoff from any MPE in the game.
@@ 3 Commitment Payoff Theorems
This section contains the main results of this paper, which characterize the payoff set in the complete information game with and without reputation, in sequential and Markovian equilibria. Informally, I establish the following table, which summarizes the takeaway from the results in this section. Together, these results form the main contribution of this paper.
@@ 3.1 Complete Information
Consider the complete information game where there is no uncertainty about the long-run player’s rationality. Say a folk theorem holds in the game if the set of payoffs attainable as part of some equilibrium in the complete-information game is the set of feasible and individually rational payoffs for the long-run player (here the set $V_{0},V^{*}]$), supposing short-run players’ myopically best reply. Proposition 1 characterizes exactly when a folk theorem holds in our game. Arguments from Fudenberg and Levine (1994), Fudenberg et al. (1994) (which would immediately imply a folk theorem) do not necessarily apply because exit can “loom large,” and thus lead to periods where the effective discount rate is small even without discounting, i.e. $\delta=1$. Thus, there exist games where not all feasible, individually rational payoffs are attainable. Proposition 1 characterizes, as a function of long-run player payoffs, exactly when a folk-like theorem holds as a function of the long-run player’s payoff, and formalizes the sense in which efficient payoffs can be sustained only when the future “looms large” given the exit structure. Let $L(a,c)=\frac{f(a,c)}{f(a,0)}$ be the likelihood ratio of exit when the long-run player chooses $c$ at some action instead of 0 given short-run player action $a$.
@@
If either of the two (equivalent) statements above holds, say the game admits a folk theorem. Proposition 1 gives a good baseline for when all feasible payoffs can be obtained in some equilibrium; moreover, the relevant incentive compatibility in Proposition 1 will give, under some monotonicity and uniqueness conditions, an exact characterization of when the long-run player exactly secures their Stackelberg payoff (Corollary 1). The converse direction of the result implies the long-run player’s best equilibrium payoff is bounded away from their Stackelberg payoff even as players get extremely patient (i.e. the folk theorem fails) so long as the incentive compatibility condition fails.
@@
Importantly, though, the equilibria constructed in the proof of Proposition 1 to sustain a high payoff for the long-run player require that punishments off the equilibrium path are non-Markov, in the sense that they are not measurable in $\mu_{t}(h^{t})$ (which is constant in the complete information case). Given this observation, it is perhaps unsurprisingly that the Markov restriction can strongly refine the set of payoffs: in the complete information game, it does so exactly by holding the long-run player down to their exact static payoff.
@@ 3.2 Guaranteeing Stackelberg Payoffs
What classes of reputation results might one hope for in the game with exit? Recall that classical reputation results hold only after many111111“Many” here depends formally on the speed of learning and the prior belief of rationality; see Gossner (2011) for a more formal discussion. (finite) periods have passed, as the long-run player must teach their opponents that they are playing the commitment action. Yet in the game with exit, they may not have the luxury of time; the game may have ended far before short-run player beliefs about long-run play have concentrated on $c^{*}$, keeping the long-run player from securing a high payoff.
Thus exit perturbs the long-run player’s strategic incentives in two ways: first, by making the future less valuable (and hence the promise of future payoffs less motivating to discipline signaling behavior today), and second, by shortening the length of time they have to convince short-run player beliefs about their (committed) play.
The first problem, as evidenced by Proposition 1, can be resolved by adding an appropriate incentive compatibility constraint for the long-run player. The second problem, however, is not so easy to fix: indeed, I will need the full force of the Markov refinement in order to iron out beliefs which are adversarial to guaranteeing the long-run player their Stackelberg payoff. The technical contribution of Theorem 1 is that these two ingredients exactly characterize when one can obtain a commitment payoff theorem in the game with exit.
@@
Say the long-run player satisfies (weak) $c^{*}$ incentive compatibility if for each $a\in\mathcal{N}(\alpha c)$ for $\alpha\in[0,1]$,
$(V_{*})L(a,c^{*})-\overline{V}_{0}>(\geq)\left(v(a,0)-v(a,c^{*})L(a,c^{*})\right)$
This incentive compatibility condition is stronger than the one found in the folk theorem, Proposition 1. Intuitively, it requires that for all potential short-run player best replies $a$ that can arise on the equilibrium path, the future looms sufficiently large that playing $c^{*}$ is enforceable against playing 0 when the long-run player can secure their Stackelberg payoff from $c^{*}$ and static Nash otherwise. In the leading application, where the likelihood ratio and stage-game payoffs depend only on $c$ (i.e. $L(a,c^{*})=L(c^{*})$ and $v(a,c)=v(c)$), this condition boils down exactly to the incentive compatibility condition in Proposition 1 (see Corollary 1). For now, it delivers the following theorem, (modulo one last definition for the converse).
@@
Primitives $(f,v,u)$ are monotone if $A$ can be ordered such that for each fixed $c$, $f(a,c)$ is strictly decreasing in $A$, $v(a,c)$ is nonincreasing, and $\mathcal{N}(\alpha c)$ is decreasing (in the strong set order) in $\alpha$.
@@
Let $\Omega=\{0,c^{*}\}$ with $\mu_{0}(\omega_{R})<1$. If the long-run player satisfies $c^{*}$ incentive compatibility, then in all Markov equilibrium $\sigma_{L}(h^{t})\equiv c^{*}$. In particular, $\underline{U}^{M}=V_{*}$.
Conversely, if $(f,v,u)$ is monotone and the long-run player does not satisfy weak $c^{*}$ incentive compatibility, then
$\limsup_{\mu_{0}(\omega_{R})\to 1}|\underline{U}^{M}-V_{*}|>0.$
@@
Let $(\sigma,\mu)$ be Markov perfect. Then for all $h^{t}$, $\text{supp}(\sigma_{L}(h^{t}))\in\{0,c^{*}\}$.
Lemma 2, along with Lemma 1 are enough to imply (following an argument similar to the first part of the proof of Proposition 1 and taking $\alpha=1$), that always playing $c^{*}$ is a Markov perfect equilibrium.
Our last step is to show that this is the only Markov perfect equilibrium. Clearly, always playing 0 cannot be Markov perfect under assumption. To see why, take $\alpha=0$ and fix any $a\in\mathcal{N}(0)$. Rearranging the incentive compatibility constraint implies
@@
so the long-run player can strictly profitably deviate by always choosing $c^{*}$ in any period they were supposed to deterministically play 0. Finally, we show that the long-run player prefers not to mix in Markov equilibria.
@@
The proof of the converse is constructive and can be found in the PROOF OF THEOREM 1 (Converse) The basic idea is that, at any $\alpha$ where the incentive compatibility constraint holds with equality, we find an equilibrium where the long-run player mixes at the time $t=0$ public history. Monotonicity of the game then ensures that at mixing histories, they underperform their commitment payoff. The condition that $\mu_{0}$ is small rules out equilibria where $\mu_{0}(\omega_{c^{*}})\approx 1$ and the long-run player can secure a very high payoff in the first period even if they play 0. ∎
The Markov refinement here disciplines short-run player learning; long-run players, in a Markov equilibrium, begin teaching short-run players about play at time $t=0$ (as opposed to needing to slowly form their reputation). Reputation thus only has a “backwards looking” component: instead of needing to teach short-run players about future play by taking some commitment action today, they need only maintain their current reputation that they will continue to do what they have already done. Theorem 1 therefore can be interpreted as saying in reputational games with exit that this backwards-looking force is the economically relevant one, i.e. the one that is not blocked by the exit structure.
Monotonicity of the game is satisfied in natural applications, including the coordination game later studied in this paper. For example, it is satisfied whenever the exit probability increases with “more aggressive’ actions $a\in A$, $v$ depends only on the long-run player’s action, and short-run players play a strictly submodular game with unique best replies (as is the case in the globalized coordination game analyzed later) in $(a,c)$ so that higher levels of long-run player actions lead to less aggressive sets of actions. Monotonicity is sufficient to imply that when the long-run player is indifferent between several actions, they cannot secure a higher payoff than their pure Stackelberg payoff. This rules out situations where the long-run player benefits from mixed Nash equilibrium. In particular, it ensures that when mixing occurs in the first period (which is used to prove the converse to Theorem 1), the long-run player must strictly underperform their commitment payoff in such an equilibrium.
In general, the incentive compatibility condition given in the theorem can be somewhat unwieldy to interpret. However, it turns out that there is a special subclass of problems in which the ability to attain the Stackelberg payoff (uniquely) is equivalent to asking if it possible to support that payoff in the complete-information game at all.

### v6

@@ 2.2 The Repeated Game
The long-run player plays the stage game in each of infinitely many periods ($t=0,1,2,\dots$), weighting each period equally (conditional on not exiting). In each period, they face a population of $N$ short-run players who take myopic best replies.
Time-$t$ histories are elements $h^{t}=(y_{s},c_{s})_{s=0}^{t-1}$ of past short-run player signals and long-run player actions. Let $H^{t}$ denote the set of time-$t$ histories, $H=\bigcup_{t}H_{t}$ the set of all histories, and $H^{\infty}=(Y\times C)^{\infty}$ the set of all infinite histories [^9] Throughout, I will sometimes drop the dependence on $Y$ when referring to a history simply as an element of $C^{\infty}$; this is meant to be read as an equivalence class of all histories with that realization of long-run player actions, for any accompanying realization of public signals $Y$.. Short-run player strategies are functions $\sigma_{i}:H\to\Delta(A_{i})$, while the (rational) long-run player’s strategy is a function $\sigma_{L}:H\to\Delta(C)$ [^10] Following Mertens et al. (2015), there is a canonical mapping from strategies $\sigma:H\to\Delta(\mathcal{X})$ and strategies $\sigma:H^{\infty}\to\Delta(\mathcal{X})$ for $\mathcal{X}\in\{C,A\}$. When it is useful (i.e. when topologizing $\sigma$), I will use these two codomains interchangeably.. Throughout, I assume players have access to a public randomization device.
Given a sequence of actions $\{a_{t},c_{t}\}_{t=0}^{\infty}$, the long-run player’s payoff is
@@
Note payoffs here are not discounted. SUPPLEMENTAL APPENDIX studies the discounted repeated game and shows the game here is the standard limit as players become patient, but with exit. However, discounting payoffs does not add anything to the analysis and, given the exit structure, would in fact substantially complicate the statement of the results without affecting the analysis. Consequently, I omit the discount rate in the main exposition, and refer to this as the case of “extreme patience” for the long-run player.
@@
and let $c^{*}$ be a maximizer in $V_{*}$ [^11] Throughout, suppose such a maximizer exists. This is true if, for example $C$ is finite or a compact topological space.. Let $V^{*}$ be the upper Stackelberg payoff, defined analogously.
@@
The long-run player can either be a rational type, $\omega_{R}\in\Omega$, who seeks to maximize their discounted payoff, or a committed type, $\omega_{c}\in\Omega$, who plays $c$ at every possible history. Short-run players hold a full-support prior belief $\mu_{0}\in\Delta(\Omega)$ over the long-run player’s type. Write $\sigma(\cdot,\omega_{c})\equiv c$ for the strategy of a committed type $\omega_{c}$.
For any strategy $\sigma$, define an associated belief system $\mu:H\to\Delta(\Omega)$, which outputs a belief about the long-run player’s type at each $h^{t}$. Together, $(\sigma,\mu)$ is an assessment. Given an assessment, let $\overline{\sigma}_{L}(h^{t})$ denote the unconditional stage-game strategy of the long-run player at $h^{t}$, averaging over $\mu(h^{t})$, and $\sigma_{L}(h^{t};\omega)$ be the stage-game strategy of a committed type $\omega_{c}$ long-run player. Let $\sigma(h^{t})$ be the rational long-run player’s strategy. Analogously, let $\overline{\mathbb{P}}^{\sigma}$ denote the probability measure over histories generated by the unconditional strategy $\sigma$, and $\mathbb{P}^{\sigma}$ the measure generated when $\omega=\omega_{R}$. Finally, let $\sigma_{L}(\cdot|h^{t})$ be the continuation strategy of the rational-type long-run player starting from some $h^{t}$.
The main results of this paper study the set of (rational) long-run player payoffs that can be attained in sequential equilibria $(\sigma,\mu)$, defined as follows.
@@
    Sequential Rationality. At every history $h^{t}$, the continuation strategy $\sigma_{L}(\cdot|h^{t})$ maximizes the long-run player’s continuation payoff given $\{\sigma_{i}(h^{t})\}_{i\in[N]}$ and $\mu(\cdot|h^{t})$.
@@
    Consistency. There exists a sequence $(\sigma^{n},\mu^{n})$, with $(\sigma^{n},\mu^{n})\to(\sigma,\mu)$ such that [^12] Because $\sigma:H\to\Delta(C)$ maps an infinite set into another, the notion of convergence is not immediately clear. Say $(\sigma^{n},\mu^{n})$ converges to $(\sigma,\mu)$ if $\sigma^{n}\to\sigma$ converges in the strong topology (see Appendix B of Luo (2025)) and $\mu^{n}\to\mu$ converges weakly.
@@
Consistency of equilibrium profiles will be useful when evaluating the value of deviations the long-run player may consider which are off the path of play (and not supported by some commitment type); absent some consistency-type condition, the value of such deviations would be hard to define in a principled way. Note the “trembles” defined here for consistency do not change the prior belief over $\Omega$ (which would then change $\mu^{n}$). A standard compactness argument shows this is without loss of generality (one could add the requirement $\mu_{0}^{n}\to\mu_{0}$ as well), but for simplicity I omit this additional notation.
Markov perfection imposes the stronger assumption that at any two histories where short-run player posterior beliefs about long-run play are the same, players take the same actions [^13] It may seem unnatural to require the long-run player take the same action when short-run player beliefs are the same. The definition of Markov perfection can be relaxed to require only that short-run players strategies are measurable in $\mu$; however, since there is a single long-run player, their strategy can then be taken to be (without loss of generality) also measurable in $\mu$, less a public randomization device.. While such a refinement may seem a-priori restrictive, I show it arises naturally when studying the effect exit has on reputational incentives (see Theorem 1).
Let $\underline{U}$ and $\overline{U}$ be the long-run player’s lowest and highest payoff from any sequential equilibrium in the game [^14] There is some potential ambiguity about whether sequential equilibrium need to exist when $C$ is infinite (see Myerson and Reny (2020) for some discussion and general fixes). However, given $\Omega$ is finite and $C$ is countable, the problem is substantially simpler—I show in Proposition 1 and Theorem 1 that sequential equilibria as defined above exist, so in particular $\underline{U},\overline{U}$ are finite.. Similarly, let $\underline{U}^{M}$ and $\overline{U}^{M}$ be the long-run player’s lowest and highest payoff from any MPE in the game.
@@ 3 Commitment Payoff Theorems
This section contains the main results of this paper, which characterize the payoff set in the complete information game with and without reputation, in sequential and Markovian equilibria. Informally, I establish the following table, which summarizes the takeaway from the results in this section.
@@ 3.1 Complete Information
Consider the complete information game where there is no uncertainty about the long-run player’s rationality. Say a folk theorem holds in the game if the set of payoffs attainable as part of some equilibrium in the complete-information game is the set of feasible and individually rational payoffs for the long-run player (here the set $V_{0},V^{*}]$), supposing short-run players’ myopically best reply. Proposition 1 characterizes exactly when a folk theorem holds in our game. Arguments from Fudenberg and Levine (1994), Fudenberg et al. (1994) (which would immediately imply a folk theorem) do not necessarily apply because exit can “loom large,” and thus lead to periods where the effective discount rate is small even without discounting, i.e. $\delta=1$. Thus, there exist games where not all feasible, individually rational payoffs are attainable. Proposition 1 characterizes, as a function of long-run player payoffs, exactly when a folk theorem holds as a function of the long-run player’s payoff, and formalizes the sense in which efficient payoffs can be sustained only when the future “looms large” given the exit structure. Let $L(a,c)=\frac{f(a,c)}{f(a,0)}$ be the likelihood ratio of survival when the long-run player chooses $c$ at some action instead of 0 given short-run player action $a$.
@@
If either of the two (equivalent) statements above holds, say the game admits a folk theorem. If the second inequality holds with equality, say the enforceability constraint binds. Proposition 1 gives a good baseline for when all feasible payoffs can be obtained in some equilibrium; moreover, the relevant incentive compatibility in Proposition 1 will give, under some monotonicity and uniqueness conditions, an exact characterization of when the long-run player exactly secures their Stackelberg payoff (Corollary 2). The converse direction of the result implies the long-run player’s best equilibrium payoff is bounded away from their Stackelberg payoff even as players get extremely patient (i.e. the folk theorem fails) so long as the incentive compatibility condition fails.
@@
Importantly, though, the equilibria constructed in the proof of Proposition 1 to sustain a high payoff for the long-run player require that punishments off the equilibrium path are non-Markov, in the sense that they are not measurable in $\mu_{t}(h^{t})$ (which is constant in the complete information case). Given this observation, it is perhaps unsurprising that the Markov restriction can strongly refine the set of payoffs: in the complete information game, it does so exactly by holding the long-run player down to their exact static payoff.
@@
3.2 Low Payoffs and Sequential Equilibrium
Before stating the main positive reputation selection result—which is stated in terms of Markov strategies—I first consider reputation effects in general sequential equilibrium. In particular, without putting additional restrictions on the survival probability function $f$, is it possible to say anything about when the long-run player can establish a reputation outside of Markov strategies? In this section, I show the answer is no, even when the “effective” discount rate is high enough that the Stackelberg action is enforceable. Formally,
Proposition 3 (Anti-Commitment.).
Fix payoffs $(u,v)$ such that there exists $c$ where $\{0\}\in\operatorname*{\arg\!\max}_{c\in C}v(a,c)$ for all $a$. Then for any $\varepsilon>0,M>0$, there exists $f$ (not necessarily satisfying Assumption 1) such that
  1. (1)
    The game admits a folk theorem and $V^{*}\geq M$.
  2. (2)
    $\limsup_{\mu_{0}(\omega_{R})\to 1}\underline{U}\leq\varepsilon$.
Note the assumption $\{0\}=\operatorname*{\arg\!\max}_{c\in C}v(a,c)$ replaces Assumption 1; as we are constructing a function $f$ for fixed $(u,v)$ and do not verify whether restrictions on $(u,v,f)$ jointly may not hold. In spirit, it is a similar dominance requirement. Note Propositions 1 and 2 remain valid without Assumption 1.
One consequence of Proposition 3 is that as $\mu_{0}(\omega_{R})\to 1$, the set of sequential equilibrium payoffs with reputation converges to the set of sequential equilibria payoffs without reputation, in stark contrast to the baseline payoff selection results of Fudenberg and Levine (1989) without exit. Note Proposition 3 is stronger than the simple statement the long-run player’s payoff can be arbitrarily low, and thus highlights the way that endogenous exit changes incentives is more subtle than a mechanic decrease in the discount rate [^15] This simpler statement can be attained by simply setting $f(a,c)$ to be arbitrarily low for all $(a,c)$. In particular, Part (1) of Proposition 3 guarantees that (a) the endogenously chosen discount rate is still large enough that the commitment action is enforceable, and (b) that repeatedly playing the commitment action gives the long-run player a high payoff. Despite this, Part (2) of Proposition 3 states that, even in the presence of reputational incentives, the long-run player’s payoff can be arbitrarily low. Indeed, Proposition 3 is a particularly strong indeterminancy result: it says for any payoffs $(u,v)$ one can find an exit probability function such that the set of equilibrium payoffs as $\mu_{0}(\omega_{R})\to 1$ is at least $\varepsilon,M]$ for any $\varepsilon,M\geq 0$.
The core force driving this stark result is that the survival probability function can vary across actions, unlike the discount rate. Consequently, one can set the survival probability at the commitment type $(a,c)$ such that $f(a,c)\approx 1$, while at the static Nash equilibrium $(a,c)\approx 0$, which is then responsible for the divergence in payoffs between the best and worst equilibrium. However, since the commitment type and Nash equilibrium are chosen endogenously, there is an additional twist: one must guarantee with the chosen $f(a,c)$ that the actions prescribed to be the commitment and equilibrium actions are in fact so. This verification—along with the explicit equilibrium—are constructed in the PROOF OF PROPOSITION 3 Of course, for the result to hold one requires that $\mu_{0}(\omega_{c^{*}})$ is not “so high” that whatever the rational player does will not affect their payoffs, and hence Proposition 3 takes $\mu_{0}(\omega_{R})\to 1$.
The equilibrium constructed in Proposition 3—which has all players choose the worst static Nash equilibrium in the first period, and the lowest Stackelberg action profile afterwards, can be adapted to immediately give the following corollary on exit probabilities, which applies for all payoffs and all exit functions.
Corollary 1.
Suppose $f(\alpha,c)$ is constant in $c$ for all $\alpha\in\mathcal{N}(0)$. Then $\underline{U}\leq\max_{\alpha\in\mathcal{N}(0)}f(\alpha,0)V_{*}$.
Corollary 1 implies that so long as the long-run player benefits from commitment, in the sense that $V_{*}>\bar{V}$, then their worst equilibrium payoff is bounded away from their Stackelberg payoff. Note Corollary 1 is distinct from mechanically lowering the discount rate, and relies on endogeneity of $f$ to player actions. If $\delta$ was simply mechanically lowered, this would be reflected in $V_{*}$, and $\underline{U}$ need not be bounded away from $V_{*}$. Indeed the equilibrium in Corollary 1 (constructed in the proof of Proposition 3) requires nonstationary play by all players, and for play to vary even when beliefs about the long-run player’s type are constant.
@@ 3.3 Guaranteeing Stackelberg Payoffs
Having shown that reputation effects no longer select payoffs in general sequential equilibria, a natural next question is whether there is a refined class of equilibria under which one can obtain a positive payoff selection result in the game with exit.
Recall that classical reputation results hold only after many161616“Many” here depends formally on the speed of learning and the prior belief of rationality; see Gossner (2011) for a more formal discussion. (finite) periods have passed, as the long-run player must teach their opponents that they are playing the commitment action. Yet in the game with exit, they may not have the luxury of time; the game may have ended far before short-run player beliefs about long-run play have concentrated on $c^{*}$, keeping the long-run player from securing a high payoff. Thus exit perturbs the long-run player’s strategic incentives in two ways: first, by making the future less valuable (and hence the promise of future payoffs less motivating to discipline signaling behavior today), and second, by shortening the length of time they have to convince short-run player beliefs about their (committed) play.
The first problem, as evidenced by Proposition 1, can be resolved by adding an appropriate incentive compatibility constraint for the long-run player. The second problem, however, is not so easy to fix: indeed, I will need the full force of a Markovian refinement in order to iron out beliefs which are adversarial to guaranteeing the long-run player their Stackelberg payoff. The technical contribution of Theorem 1 is that these two ingredients exactly characterize when one can obtain a commitment payoff theorem in the game with exit.
@@
Say the long-run player satisfies (weak) $c^{*}$ incentive compatibility if for each $\alpha\in\mathcal{N}(\beta c^{*})$ and for all $\beta\in[0,1]$,
$(V_{*})L(\alpha,c^{*})-\overline{V}_{0}>(\geq)\left(v(\alpha,0)-v(\alpha,c^{*})L(\alpha,c^{*})\right)$
Here I abuse notation and use $\beta c$ to refer to the lottery where the long-run player chooses $c$ with probability $\beta$ and 0 with probability $1-\beta$.
This incentive compatibility condition is a little stronger than the one found in the folk theorem, Proposition 1. Intuitively, it requires that for all potential short-run player best replies $a$ that can arise on the equilibrium path, the future looms sufficiently large that playing $c^{*}$ is enforceable against playing 0 when the long-run player can secure their Stackelberg payoff from $c^{*}$ and static Nash otherwise. In particular, it requires that no matter how short-run players break ties, $c^{*}$ is enforceable—unlike the incentive compatibility condition in Proposition 1, which assumes ties are broken in favor of the long-run player. Yet there are settings where the two conditions coincide; in the leading application, where the likelihood ratio and stage-game payoffs depend only on $c$ (i.e. $L(a,c^{*})=L(c^{*})$ and $v(a,c)=v(c)$), this condition boils down exactly to the incentive compatibility condition in Proposition 1 (see Corollary 2). For now, it delivers the following theorem, (modulo one last definition for the converse).
@@
Primitives $(f,v,u)$ are monotone if $A$ can be ordered such that for each fixed $c$, $f(a,c)$ is strictly decreasing in $A$, $v(a,c)$ is nonincreasing, and $\mathcal{N}(\beta c)$ is decreasing (in the strong set order) in $\beta$.
@@
Fix $\Omega=\{\omega_{R},\omega_{c^{*}}\}$ and $\mu_{0}(\omega_{R})<1$ . If the long-run player satisfies $c^{*}$ incentive compatibility, then in all Markov equilibrium $\sigma_{L}(h^{t})\equiv c^{*}$. In particular, $\underline{U}^{M}=V_{*}$.
Conversely, if $(f,v,u)$ is monotone and the long-run player does not satisfy weak $c^{*}$ incentive compatibility, then $\limsup_{\mu_{0}(\omega_{R})\to 1}|\underline{U}^{M}-V_{*}|>0$.
Several features of Theorem 1 are worth discussing. First is the role of $c^{*}$-incentive compatibility, which ensures both that $c^{*}$ is enforceable (when $\beta=1$) and that it remains enforceable when short-run players best respond to a mixture of 0 and $c^{*}$. The latter is critical to discipline long-run player incentives and ensure mixed-strategy Markov equilibria do not arise on the path of play (see Lemma 3).
The second is the role of the Markov refinement, which is best understood in comparison to Proposition 3. In particular, the Markov refinement disciplines the feasible payoff set after rationality is revealed (thus decreasing incentives for “revealing rationality”), making it hard for short-run players to reward in the future actions which are non-Stackelberg. It also ensures short-run players believe the commitment action will be played whenever reputation foundations have bite, which ensures that exit does not “loom large” in the first period. Markov beliefs thus rule out (potentially pathological) equilibria where the long-run player is rewarded in the future in a complicated way for revealing they are rational today, even though doing so will cause the game to end with arbitrarily high probability today.
The Markov refinement can thus be more broadly interpreted as follows. In general reputation games, there are two forces at play: a backwards looking one (is past long-run player behavior consistent with play by the commitment type?) and a forward looking one (does this imply future long-run player behavior will be consistent with Stackelberg play?) In traditional reputation models, these two forces together ensure the long-run player can eventually teach short-run players their action is indistinguishable from that of the committed type regardless of the equilibria, guaranteeing a high payoff bound. The Markov refinement preserves the backwards looking force (learning starts at $t=0$), but cannot preserve the forward looking one since exit might occur before “future long-run player behavior” matters. Consequently, the Markov refinement shows that in games with exit, measurability of play in beliefs is a substitute to the forward-looking force, and preserves a payoff selection theorem.
Finally, the converse. Monotonicity of the game is satisfied in natural applications, including the coordination game later studied in this paper. For example, it is satisfied whenever the exit probability increases with “more aggressive’ actions $a\in A$, $v$ depends only on the long-run player’s action, and short-run players play a strictly submodular game with unique best replies (as is the case in the globalized coordination game analyzed later) in $(a,c)$ so that higher levels of long-run player actions lead to less aggressive sets of actions. Monotonicity is sufficient to imply that when the long-run player is indifferent between several actions, they cannot secure a higher payoff than their pure Stackelberg payoff. This rules out situations where the long-run player benefits from mixed Nash equilibrium. In particular, it ensures that when mixing occurs in the first period (which is used to prove the converse to Theorem 1), the long-run player must strictly underperform their commitment payoff in such an equilibrium. I next give an outline of the proof of Theorem 1.
@@
Let $(\sigma,\mu)$ be Markov perfect. Then for all $h^{t}$, $\text{supp}(\sigma_{L}(h^{t}))\subset\{0,c^{*}\}$.
Lemma 2, along with Lemma 1 are enough to imply (following an argument similar to the first part of the proof of Proposition 1 and taking $\beta=1$), that always playing $c^{*}$ is a Markov perfect equilibrium.
Our last step is to show that this is the only Markov perfect equilibrium. Clearly, always playing 0 cannot be Markov perfect under assumption. To see why, take $\beta=0$ and fix any $a\in\mathcal{N}(0)$. Rearranging the incentive compatibility constraint implies
@@
so the long-run player can strictly profitably deviate by always choosing $c^{*}$ in any period they were supposed to deterministically play 0 where $\mu(h^{t})(\omega_{R})<1$, e.g. the first period. Finally, we show that the long-run player prefers not to mix in Markov equilibria.
@@
The proof of the converse is constructive and can be found in the PROOF OF THEOREM 1 (Converse) The basic idea is that, at any $\beta$ where the incentive compatibility constraint holds with equality, we find an equilibrium where the long-run player mixes at the time $t=0$ public history. Monotonicity of the game then ensures that at mixing histories, they underperform their commitment payoff. The condition that $\mu_{0}$ is small rules out equilibria where $\mu_{0}(\omega_{c^{*}})\approx 1$ and the long-run player can secure a very high payoff in the first period even if they play 0. ∎
In general, the incentive compatibility condition given in the theorem can be somewhat unwieldy to interpret. However, it turns out that there is a special subclass of problems in which the ability to attain the Stackelberg payoff (uniquely) is equivalent to asking if it is possible to support that payoff in the complete-information game at all.